OpenStudy (anonymous):

explain why the commutative property does not hold true for subtraction or division

OpenStudy (anonymous):

anyone please?

OpenStudy (anonymous):

because the commutative law for addition says \(a+b=b+a\) like for example \[10+5=5+10\]but if you replace \(+\) by \(-\) it is not true that \(a-b=b-a\) like for example \[10-5\neq 5-10\]